What's the linear equation for '5y - 2x + 7 =0' and m= ?, b=?
Answer by rapaljer(4671)

Solve for y to write this in the form y=mx+b.

5y - 2x + 7 =0
5y = 2x-7
y=1/5x -7/5

slope m= 1/5
y-intercept b= -7/5

R^2

For more help with the topic of graphing and slope of a line, please see my own website. Do a "Bing" search for my last name "Rapalje", and look for "Rapalje Homepage" near the top of the Bing Search. Near the top of my Homepage is a link "Basic, Intermediate and College Algebra: One Step at a Time." Click on this link and choose "Basic Algebra". Look for "Chapter 4: Graphing". Several sections are there with a "non-traditional explanation that my students told me is easier to understand than the traditional textbooks. Also, look for my "MATH IN LIVING COLOR" pages for most of these sections, and also a video of me teaching my class. To see the video (which is FREE!), look on my Homepage for the link "Rapalje Videos in Living Color", choose "Basic Algebra", and select "Graphing and Slope of a Line."

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus